Oct. 3   Purchased merchandise from Hoagie Co., $9,950, terms FOB shipping point, 2/10, n/eom. Prepaid freight of $220 was added to the invoice.

Debit Credit

Oct. 13   Paid Hoagie Co. for invoice of October 3.

Debit Credit
Answer #1
Under Gross Method:
Debit Credit
Oct.3 Merchandise Inventory 10170 =9950+220
         Accounts Payable-Hoagie Co. 10170
Oct.13 Accounts Payable-Hoagie Co. 10170
         Merchandise Inventory 199 =9950*2%
        Cash 9971
Under Net Method:
Debit Credit
Oct.3 Merchandise Inventory 9971 =9950*(1-0.02)+220
         Accounts Payable-Hoagie Co. 9971
Oct.13 Accounts Payable-Hoagie Co. 9971
        Cash 9971
Note: Solution under both gross and net method have been provided, the appropriate solution
can be used based on instructions provided in the question